OpenStudy (anonymous):

Solve sin A + sin 2A + sin 3A + sin 4A = 0, for 0 < A < 180 (0 and 180 inclusive)

OpenStudy (anonymous):

I've tried using factor formula but still did not manage to get the answer, not sure if factor formula is the right method

OpenStudy (anonymous):

I rearrange to sin 4A + sin 2A + sin 3A + sin A = 0, and after applying factor formula, 2 sin 3A cos A + 2 sin 2A cos A = 0 2 cos A ( sin 3A + sin 2A) = 0 2 cos A ( sin 5/2 A cos 1/2 A) = 0

OpenStudy (anonymous):

@thomaster @Preetha

OpenStudy (mathmath333):

\(A=0^{\circ}\) is one solution by obsrvation

OpenStudy (welshfella):

@HatcrewS you are on the right road to solving it

OpenStudy (anonymous):

@welshfella how do I continue from there ? T.T

OpenStudy (welshfella):

i think there is an error in your last line

OpenStudy (welshfella):

2 cos A ( sin 5/2 A cos 1/2 A) = 0 should be 2 cos A ( 2 sin 5/2 A cos 1/2 A) = 0 now its just a matter of equating the 3 terms to 0

OpenStudy (anonymous):

@welshfella oh okay i'll try again thanks!

OpenStudy (welshfella):

2 cos A = 0 A = 90 degrees

OpenStudy (anonymous):

I solve and got 0, 90, 144, 180 but im still missing 72(answer sheet shows that 72 is one of the answer). Can someone help me get 72 ?